OpenStudy (anonymous):

Find the height of a triangle with area of 81 square centimeters and a base of 9 cm. Find the height of a triangle with area of 81 square centimeters and a base of 9 cm. @Mathematics

OpenStudy (anonymous):

what is the formula of area of triangle?

OpenStudy (yash2651995):

ar = (1/2)x base x ht

OpenStudy (anonymous):

thanks yash

OpenStudy (anonymous):

magar jisse pucha tha wahi gayab ho gaya

OpenStudy (anonymous):

wanted to get it solved from her only

OpenStudy (yash2651995):

delete kar dun comment??

OpenStudy (anonymous):

yeah kiss69fist23 can u tell me the height of this triangle

OpenStudy (anonymous):

no i dont kno Mujhe samajha mem nahim ata

OpenStudy (anonymous):

hahahahaaha u too indian

OpenStudy (anonymous):

:D haha yes

OpenStudy (anonymous):

whr from

OpenStudy (anonymous):

see the formula for Area of triangle is as yash has written above

OpenStudy (anonymous):

i dint see it :(

OpenStudy (anonymous):

can u guess where im from haha

OpenStudy (yash2651995):

"area = (1/2)x base x perpendicular ht from that base"

OpenStudy (anonymous):

yeah sure i can guess u must be from somewhere in india only

OpenStudy (anonymous):

hahaha of course dear sheg

OpenStudy (anonymous):

Area of triangle = 1/2(base)*(height)

OpenStudy (anonymous):

Sukriya

OpenStudy (yash2651995):

hey am i invisible??

OpenStudy (anonymous):

LOL my Apologies :D THAnk youe dearest yash :D

OpenStudy (anonymous):

ok now can u solve it

OpenStudy (anonymous):

got it :)

OpenStudy (anonymous):

81sq cm = 1/2 (9cm)(height) Height = 162/9 = 18 cm